OpenStudy (anonymous):

645 = 5x(to the 3rd power) solve for x how to solve this??

OpenStudy (aum):

\(\large 645 = 5x^3\) or \(\large 645 = (5x)^3\) ?

OpenStudy (anonymous):

yea first one

OpenStudy (aum):

To solve for x, first isolate x. To do that divide both sides by 5 Then take the cube root on both sides and that will give you x.

OpenStudy (anonymous):

also how do you raise the 3?

OpenStudy (aum):

you mean how to type 3 as an exponent so it appears raised on this website?

OpenStudy (aum):

What do you mean by: also how do you raise the 3?

OpenStudy (anonymous):

yeah how to type an exponent.. so the answer is 129?

OpenStudy (aum):

No. The answer is cube root of 129. That is, \(\large x= \sqrt[3]{129} \). To type exponents, radical sign, etc., use the Equation Editor. Just click on the blue button marked Equation below the reply box and explore various things you can do with an equation editor.

OpenStudy (anonymous):

my hw reads to write answer to the hundredths place, so if i put \[3\sqrt{129}\] should i put 34.07?